# Robust Permissive Controller Synthesis for Interval MDPs

### v1

@@ Abstract
We address the problem of robust permissive controller synthesis for robots operating under uncertain dynamics, modeled as Interval Markov Decision Processes (IMDPs). IMDPs generalize standard MDPs by allowing transition probabilities to vary within intervals, capturing epistemic uncertainty from sensing noise, actuation imprecision, and coarse system abstractions—common in robotics. Traditional controller synthesis typically yields a single deterministic strategy, limiting adaptability. In contrast, permissive controllers (multi-strategies) allow multiple actions per state, enabling runtime flexibility and resilience. However, prior work on permissive controller synthesis generally assumes exact transition probabilities, which is unrealistic in many robotic applications. We present the first framework for robust permissive controller synthesis on IMDPs, guaranteeing that all strategies compliant with the synthesized multi-strategy satisfy reachability or reward-based specifications under all admissible transitions. We formulate the problem as mixed-integer linear programs (MILPs) and propose two encodings: a baseline vertex-enumeration method and a scalable dualization-based method that avoids explicit enumeration. Experiments on four benchmark domains show that both methods synthesize robust, maximally permissive controllers and scale to large IMDPs with up to hundreds of thousands of states.
@@ I Introduction
Decision-making under uncertainty is a central challenge in robotics [1, 2]. Robots operate in environments where sensing is noisy, actuation is imprecise, and system abstractions are inevitably coarse. Standard Markov decision processes (MDPs) are widely used to model sequential decision-making, but they assume exact transition probabilities, which are rarely available in practice. Recent surveys highlight the broad landscape of uncertainty models [3, 4].
@@
Among these models, Interval MDPs (IMDPs) play a critical role. By allowing transition probabilities to vary within intervals, IMDPs capture epistemic uncertainty while retaining a tractable structure for analysis and synthesis. Robust controller synthesis for IMDPs has been studied extensively, with applications ranging from robotic planning and verification to hybrid and neural network dynamics [5, 6, 7, 8]. However, existing methods typically return a single policy—often deterministic or randomized—but not permissive, which can be overly restrictive in practice.
@@
In contrast, permissive controllers, also known as multi-strategies, allow multiple actions to remain enabled in each state. Such controllers preserve flexibility at runtime: a robot may adaptively select among admissible actions based on external conditions, unforeseen disturbances, or higher-level objectives, while still maintaining correctness. Permissive synthesis has been studied for MDPs and probabilistic games with exact transitions, often via mixed-integer linear programming (MILP) encodings that maximize permissiveness subject to property satisfaction [9, 10, 11, 12]. Yet, existing permissive methods assume precise probability models and do not account for epistemic uncertainty as captured by IMDPs.
The lack of robust permissive synthesis methods creates a significant gap for robotics applications. Consider a mobile robot navigating in an uncertain environment. A deterministic controller may commit to a single action sequence that is safe under all transition realizations, but in doing so, it may forgo many other safe alternatives. A permissive controller, in contrast, enables all actions consistent with safety, allowing adaptation at runtime. To be useful in robotics, such permissive strategies must be robust to transition uncertainty, guaranteeing correctness under all admissible distributions of the IMDP.
Another way to interpret permissive controllers is as shields for safe reinforcement learning (RL) [12, 13], which filter unsafe actions online while leaving other choices unrestricted. A permissive multi-strategy serves precisely this role, specifying admissible actions in each state and blocking the rest. Prior permissive synthesis methods thus align with shields in known MDPs, while robust IMDP synthesis aligns with worst-case reasoning under uncertainty. Bridging these ideas, robust permissive synthesis over IMDPs naturally yields robust shields: multi-strategies that guarantee safety against all admissible transitions while minimally restricting learning or heuristic policies. This view is particularly relevant for emerging paradigms such as meta-RL [14] and in-context RL [15], where agents must adapt across distributions of MDPs. Both are attracting interest in robotics, but ensuring safety under distributional variability remains challenging; robust permissive shields offer a principled way to address this while preserving flexibility for adaptation.
@@
In this paper, we present the first framework for robust permissive controller synthesis on IMDPs. We focus on reachability and reward-based specifications, which are widely used in robotics. Our approach guarantees that every strategy compliant with the synthesized multi-strategy satisfies the specification under all admissible transitions. We formulate the synthesis problem as mixed-integer linear programs and propose two alternative encodings:
@@
    Dualization-based encoding, which avoids explicit enumeration by leveraging robust optimization, yielding a compact formulation that scales linearly with the number of successors.
We implement both encodings in a prototype tool and evaluate them on four representative benchmark domains: aircraft collision avoidance, semi-autonomous vehicle control, obstacle navigation, and warehouse navigation. Our experiments show that both methods synthesize robust, maximally permissive controllers, scaling to IMDPs with hundreds of thousands of states.
By bridging permissive multi-strategy synthesis with robust IMDP reasoning, our work advances the state of the art in uncertainty-aware decision-making for robotics. It provides both theoretical foundations and practical algorithms for synthesizing controllers that are robust to model uncertainty while retaining runtime flexibility.
@@ II Related Work
Permissive Controller Synthesis. Permissive controller synthesis was introduced to preserve flexibility by allowing multiple actions per state while still satisfying specifications. Initial work encoded permissive multi-strategy synthesis as a MILP with penalties for disabled actions [9]. Subsequent studies extended this idea to multi-objective settings, synthesizing permissive controllers that accommodate uncertain human preferences [11], and to robust planning in models with set-valued transitions under temporal-logic objectives [31]. Beyond formal synthesis, permissive controllers are closely related to shields in safe reinforcement learning, which block unsafe actions online while leaving other choices unrestricted; in this setting, a permissive multi-strategy naturally serves as a shield that enforces safety with minimal interference [12, 13]. While these approaches demonstrate the utility of permissive strategies, they assume exact or set-valued probabilities. To the best of our knowledge, no prior work addresses robust permissive synthesis for IMDPs under interval uncertainty, which is the gap filled by this paper.
@@ V-C Results Analysis
The results in Table I show that both encodings are capable of synthesizing maximally permissive robust strategies. In all benchmark instances, the normalized permissiveness is close to 1, confirming that the synthesized strategies preserve nearly all available actions.
@@ VI Conclusion
We presented the first framework for robust permissive controller synthesis on IMDPs, enabling controllers that guarantee specification satisfaction under all admissible transitions while preserving multiple action choices for runtime flexibility. The problem was formulated as mixed-integer linear programs with two encodings: a conceptually simple vertex-enumeration method and a dualization-based method that scales linearly with the number of successors. Experiments on four benchmark domains showed that both approaches synthesize maximally permissive robust strategies, with the dualization method offering superior scalability on large models.

### v2

# Optimization-Based Robust Permissive Synthesis for Interval MDPs

@@ Abstract
We present an optimization-based framework for robust permissive synthesis for Interval Markov Decision Processes (IMDPs), motivated by robotic decision-making under transition uncertainty. In many robotic systems, model inaccuracies and sensing noise lead to interval-valued transition probabilities. While robust IMDP synthesis typically yields a single policy and permissive synthesis assumes exact models, we show that robust permissive synthesis under interval uncertainty can be cast as a global mixed-integer linear program (MILP) that directly encodes robust Bellman constraints. The formulation maximizes a quantitative permissiveness metric (the number of enabled state–action pairs), while guaranteeing that every compliant strategy satisfies probabilistic reachability or expected reward specifications under all admissible transition realizations. To address the exponential complexity of vertex-based uncertainty representations, we derive a dualization-based encoding that eliminates explicit vertex enumeration and scales linearly with the number of successors. Experimental evaluation on four representative robotic benchmark domains demonstrates scalability to IMDPs with hundreds of thousands of states. The proposed framework provides a practical and general foundation for uncertainty-aware, flexibility-preserving controller synthesis in robotic systems.
@@ I Introduction
Decision-making under uncertainty is a fundamental challenge in robotics [35, 20]. Robots operate under sensing noise, actuation imprecision, and modeling errors, yet standard Markov decision processes (MDPs) assume exact transition probabilities. Among uncertainty models, Interval MDPs (IMDPs) capture epistemic uncertainty by allowing transition probabilities to vary within intervals, while remaining amenable to formal analysis and synthesis [4, 33]. Robust controller synthesis for IMDPs has been widely studied in robotic planning, verification, and learning-based control [17, 18, 8, 1]. However, existing methods typically synthesize a single policy—deterministic or randomized—which can be overly restrictive in practice.
@@
In contrast, permissive controllers (multi-strategies) retain multiple admissible actions per state, preserving flexibility at runtime while maintaining correctness. Permissive synthesis has been developed for MDPs and probabilistic games with exact transition models, often via mixed-integer linear programming formulations that maximize permissiveness subject to specification satisfaction [11, 19, 9, 16]. Yet these approaches assume precise transition probabilities and do not address epistemic uncertainty as captured by IMDPs. For robotics applications, permissiveness must therefore be combined with robustness, ensuring correctness under all admissible transition realizations.
Permissive controllers can also be viewed as shields for safe reinforcement learning (RL) [16, 2], filtering unsafe actions online while leaving other choices unrestricted. Recent work has used IMDP abstractions to construct data-driven shields for unknown continuous systems [30], highlighting the promise of interval-based models for safety-critical learning. Robust permissive synthesis over IMDPs naturally unifies these ideas, yielding shields that provide worst-case safety guarantees while minimally restricting adaptive or learning-based policies. This perspective is particularly relevant for emerging paradigms such as meta-RL [5] and in-context RL [25], where agents must operate across varying environments while maintaining safety.
@@
In this paper, we present a framework for robust permissive controller synthesis on IMDPs. We consider probabilistic reachability and expected reward specifications, which are widely used in robotics. Our approach guarantees that every strategy compliant with the synthesized multi-strategy satisfies the specification under all admissible transition probabilities. We formulate the synthesis problem as mixed-integer linear programs and develop two alternative encodings:
@@
    Dualization-based encoding, which avoids explicit vertex enumeration via robust optimization, yielding a compact formulation that scales linearly with the number of successors.
We implement both encodings in a prototype tool and evaluate them on four benchmark domains: aircraft collision avoidance, semi-autonomous vehicle control, obstacle navigation, and warehouse navigation. The results demonstrate that both approaches synthesize robust and highly permissive controllers, scaling to IMDPs with hundreds of thousands of states. By unifying permissive multi-strategy synthesis with robust IMDP reasoning, this work advances uncertainty-aware decision-making for robotics. It provides both theoretical guarantees and practical algorithms for synthesizing controllers that remain correct under model uncertainty while preserving runtime flexibility.
@@ II Related Work
Permissive Controller Synthesis. Permissive controller synthesis was introduced to preserve flexibility by allowing multiple actions per state while still satisfying specifications. Initial work encoded permissive multi-strategy synthesis as a MILP with penalties for disabled actions [11]. Subsequent studies extended this idea to multi-objective settings, synthesizing permissive controllers that accommodate uncertain human preferences [9], and to robust planning in models with set-valued transitions under temporal-logic objectives [38]. Beyond formal synthesis, permissive controllers are closely related to shields in safe reinforcement learning, which block unsafe actions online while leaving other choices unrestricted; in this setting, a permissive multi-strategy naturally serves as a shield that enforces safety with minimal interference [16, 2]. These approaches, however, assume exact or set-valued transition models and do not address interval uncertainty.
Recent work has extended shielding ideas to IMDP abstractions of unknown continuous systems [30], where unsafe actions are iteratively pruned using value iteration until a worst-case violation probability threshold is satisfied. The resulting shield is inclusion-maximal with respect to the safety bound and focuses on safe LTL specifications. In contrast, we formulate robust permissive synthesis for IMDPs as a single global optimization problem. Our approach encodes robust Bellman constraints directly in a mixed-integer linear program and maximizes a permissiveness metric (the total number of enabled state–action pairs), while supporting probabilistic reachability and expected reward objectives.
@@ V-B Experimental Setup
Our evaluation focuses on assessing the scalability and optimality of the proposed MILP encodings; direct comparison with IMDP-based shielding approaches such as [30] is outside our scope due to differences in problem formulation and objectives.
@@ V-C Results Analysis
The results in Table I show that both encodings successfully synthesize maximally permissive robust strategies. In all benchmark instances, the normalized permissiveness is close to 1, confirming that the synthesized strategies preserve nearly all available actions.
@@ VI Conclusion
We presented a framework for robust permissive controller synthesis on IMDPs, enabling controllers that guarantee specification satisfaction under all admissible transitions while preserving multiple action choices for runtime flexibility. The problem was formulated as mixed-integer linear programs with two encodings: a conceptually simple vertex-enumeration method and a dualization-based method that scales linearly with the number of successors. Experiments on four benchmark domains showed that both approaches synthesize maximally permissive robust strategies, with the dualization method offering superior scalability on large models.